40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jul 26 11:21:53 2016, awade and Antonio, DailyProgress, , Initial search for beat note refcavs' beatnote 
    Reply  Tue Jul 26 12:56:13 2016, Koji, DailyProgress, , Initial search for beat note refcavs' beatnote 
       Reply  Tue Jul 26 15:12:53 2016, awade, DailyProgress, BEAT, Initial search for beat note refcavs' beatnote 
          Reply  Wed Jul 27 23:05:45 2016, awade and Antonio, DailyProgress, BEAT, Initial search for beat note refcavs' beatnote 
             Reply  Fri Jul 29 16:29:53 2016, awade, DailyProgress, BEAT, Initial search for beat note refcavs' beatnote 20160727_DirectLaserBeatNoteDetector.pdf
Message ID: 1684     Entry time: Tue Jul 26 11:21:53 2016     Reply to this: 1687
Author: awade and Antonio 
Type: DailyProgress 
Category:  
Subject: Initial search for beat note refcavs' beatnote 

With cavities mode matched and PDH locked on reflection we are now looking to find the beat notes between the two separate cavity-laser systems from the transmitted refcavity beams.

The transmitted beams were realigned through the combining beam splitter using the existing iris positions.  It appears that the the beam alignments, after various adjustments, were coincident on the 125 MHZ beat note detector but at different angles.  One would expect to see a beat note but reduced by the fringe across the beam front. As it is best practice to have overlapping and co-propagating beams, the south path and then the north path were walked to correspond with iris located on the unused port of the PLL (phase locked loop) combining beam splitter. 

We searched for a beat note over a selection of laser temperature offsets.  The South laser was locked to its refcavity at 0.8,1.68,2.53, 3.50 and 4.36 V slow control offset* and then the North cavity was locked at all combinations of 0.475, 1.359,2.252,3.147,4.129,5.042,5.965,6.976 V.  We did not observe a beat note when looking at the PD RF output on a spectrum analyzer. It is very possible that the corresponding FSRs of the two cavities just were not within the bandwidth of the detectors.  We should look around for a RF detector with >=1 GHz bandwidth and place at a pick off somewhere further up the laser path as a way of quickly checking the laser offset without the need to also align the cavity resonance.  This would save us a lot of time and pain carefully walking the refcavity temperature until we had corresponding overlap.  Once we have known operating point this process will be much easier, but a second RF PD would be a very useful diagnostic tool to have; we should generally try to do things the easy way.

 

*Note that there is a mode hopping region towards the top of the south cavity's range wich prevents going any higher in temperature offset (using the front panel voltage) for now.

ELOG V3.1.3-